Can someone help me please?
The average number of diners in a restaurant can be modeled with the function
D(t)= -2t^2+20t+15 , where t is the number of hours after 8:00 A.M. At what time is the number of diners at its maximum?
thanks so much!

Answer by stanbon(48546)

The average number of diners in a restaurant can be modeled with the function
D(t)= -2t^2+20t+15 , where t is the number of hours after 8:00 A.M. At what time is the number of diners at its maximum?
---------
You have a quadratic with a=-2 , b=20
Max occurs when t=-b/2a = -20/(2*-2)= 5 hurs
5hrs after 8:00AM is 1:PM
-----------
Cheers,
Stan H.